# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 3.8.4 Reconciling Emergences
Based on the concept of synergistic information in the PID framework, Rosas proposed a quantitative definition of causal emergence and further classified it using the $\phi$ID framework, which addresses the problem of finding a coarse-graining strategy. The definition includes two aspects: the first is whether the system has the ability to produce causal emergence objectively. Then, on this basis, whether there is causal emergence happening with specific macroscopic features.
In terms of the capacity to generate causal emergence for a system, this definition links causal emergence to synergistic relationships among variables in different times. Therefore, a system $X_{t}$ is said to have capacity for causally emergent features if and only if that:
@@
This means that causal emergence is regarded as the synergy effect among variables between the previous and subsequent moments in the system with Markovian dynamics.
Then, Rosas further classify causal emergence by the $\phi$ID framework. Among the sixteen $\phi$ID atoms obtained by decomposing the mutual information $I(X_{t};X_{t+1})$ by the $\phi$ID, here are four information atoms corresponding to synergistic effect, which is regarded as causal emergence, are $I^{\{12\}\rightarrow\alpha}_{\partial}(X_{t},X_{t+1})$, $\alpha\in\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$. Since $\phi$ID includes forward and backward PID, he further divides causal emergence into downward causality and causal decoupling according to the different characteristics of those information atoms. The downward causality is defined by $\mathcal{D}(X_{t},X_{t+1})$, which encapsulating the information atoms that manifest a synergistic effect exclusively in the forward PID:
$\mathcal{D}(X_{t},X_{t+1}):=\sum_{\begin{subarray}{c}\alpha\in A\end{subarray}}{I^{\{12\}\rightarrow\alpha}_{\partial}}(X_{t},X_{t+1})$ (32)
Conversely, the causal decoupling is denoted as $\mathcal{G}(X_{t},X_{t+1})$, pinpointing the specific information atom where both forward and backward PID exhibit synergy:
@@
As the $\phi$ID is a non-overlapping decomposition of all information, this classification takes into account all cases of causal emergence, that
@@
Figure 12: Causal decoupling ($\mathcal{G}$) and downward causation ($\mathcal{D}$) represented in double redundancy lattice.
Rosas also gives out a version of causal emergence quantification depending on a macro-variable, i.e., a coarse-graining strategy. If a system has the capacity to generate causal emergence, then there can be some macroscopic features that exhibit causal emergence. A feature variable $V$ is said to be supervenient on the underlying system if it does not provide any predictive power for future states at times $t+1$ if the complete state $X$ of the system at time $t$ is known with perfect precision, which is equivalent to $V_{t}$ is statistically independent of $X_{t+1}$ when $X_{t}$ is given. Then, for a system described by $X_{t}$, a supervenient feature $V_{t}$ is said to exhibit causal causation if:
@@
For this definition, the system’s capacity of causal emergence is required, where $Syn(X_{t};X_{t+1})>0$, since $Un(V_{t};X_{t+1}|X_{t})\leq Syn(X_{t};X_{t+1})$ holds for any supervenient feature $V_{t}$. Corresponding to the classification of system’s capacity, the Downward Causation exists of a feature variable $V$ when:
@@
and the Causal Decoupling exists when
@@
which also needs to be premised on the capacity of the system. Furthermore, $V_{t}$ is said to have pure causal decoupling if $Un(V_{t};X_{t+1}^{\alpha}|X_{t})=0$ and $Un(V_{t};X_{t+1}|X_{t})>0$. If all the emergent features exhibit pure causal decoupling, then the system is said to be perfectly decoupled.
@@
Although a rigorous quantitative definition for causal emergence is proposed, the mathematical formulations used in $\phi$ID can be complex and computationally demanding, making it difficult to apply the method to real-world systems. Also, there are many calculation methods in the PID framework, which makes the quantitative definition determined by the specific quantitative method when applied. Therefore, to solve those problems, Rosas also relaxed the calculation methods of causal emergence, thus obtaining sufficient conditions for the two phenomena as identification criteria.
Specifically, in order to avoid exploring the specific quantitative relationship of synergistic and redundant information, this criteria repeatedly subtracts redundant information to make positive results a sufficient condition for emergent phenomena. Although this makes the quantitative method lose some of its generalizability, it improves the reliability of the results. The three indicators to be used are:
@@
which measures the mutual information between macro-variables across two time steps minus the information between micro-state to macro-state, where $V$ is a predefined macro-variable, but how to find such macro-variable is never discussed in Rosas et al. (2020).
@@
which is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$ in next step minus all the summation of mutual information between $X_{t}^{i}$ and $X_{t+1}^{j}$.
@@
this is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$
@@
The specific functions of the indicators are as follows: $\Psi_{t,t+1}(V)>0$ is a sufficient condition for the causal emergence of $V_{t}$. Similarly, $\Delta_{t,t+1}(V)>0$ is as sufficient condition for $V_{t}$ to show downward causation. Finally, $\Psi_{t,t+1}(V)>0$ and $\Gamma_{t,t+1}(V)=0$ is a sufficient condition for causal decoupling to occur.
@@
In summary, Rosas proposes an approach to quantitatively describe causal emergence by establishing the relationship between causal emergence and the synergistic effect of variables in different times, and further classified causal emergence by $\phi$ID. The definition not only objectively judges whether the system has the capacity to produce causal emergence, but can also measure the characteristics of causal emergence of specific macro features. The most notable contributions of his work were to bring the study of emergence closer to quantitative and empirical scientific research, to classify types of causal emergence, and to complement philosophical discussions about emergence.
There are some applications of PID or $\phi$ID. Recently, Luppi et al. apply $\phi$ID method on brain BOLD signals Luppi et al. (2022) to identify the synergistic core of the brain. They found synergistic information is for the integration of different parts to a whole, and redundancy is for robustness.
With the development of partial information decomposition techniques, we can further decompose the mutual information between two variables. This may help us to reveal the properties of systems with more aspects. For example, in  Varley and Hoel (2022), the authors first decompose the mutual information of a system, and calculate an indicator of synergy bias according to partial information decomposition method Williams and Beer (2010) to characterize how synergistic information distribute on different levels of the system. Higher synergy bias indicates that more partial information is in synergistic relationships. Then, the authors found that in some systems with occurrence of causal emergence, if the system is coarsen to a reduced system, the synergy bias will increase. Therefore, partial information is converted from redundancy to synergy as we coarse-grain the system. The conclusion is emergence is a kind of information conversion.
@@ 3.8.5 Comparison with Hoel’s framework
When comparing Hoel’s framework to Rosas’ quantification framework of causal emergence, several clear advantages can be observed in the latter. Firstly, Rosas’ theory does not require a predetermined coarse-graining method, making it more mathematically rigorous and formal. Secondly, it offers a detailed decomposition of causal emergence, specifically downward causation and causal decoupling. Lastly, it effectively avoids cases of fake causal emergence, where the macro-variable depends solely on unique or redundant information from the micro-variables.
@@
However, there are also some disadvantages to consider. Firstly, in order to obtain the full information lattice, a systematic iteration of all variable compositions is required. Additionally, despite the use of formula 35, it is necessary to define a macro-variable. Unfortunately, the authors do not provide any method to identify such a variable. Secondly, all the mutual information and its decomposition are based on correlations rather than causality. It is crucial to discuss how to incorporate causal elements, such as intervention and counterfactuals, into the framework.
@@
Finally, the issue of identifying whether causal emergence occurs in a system based on the given time series data of its behaviors has not been addressed in the preceding discussion. To tackle this problem, the application of emerging technologies in machine learning and artificial intelligence is required. These technologies can provide valuable tools and techniques for detecting and analyzing causal emergence.
4 Causal Emergence and Artificial Intelligence
@@
Recently, newly emerging artificial intelligence technologies have made significant breakthroughs in tackling a range of important and challenging problems. Examples include defeating human champions in complex games like Go Silver et al. (2016); Dabney et al. (2020), and predicting the intricate structures of protein folding Senior et al. (2020). These achievements have been made possible through the application of machine learning methods, which leverages well-designed neural network architectures and automatic differentiation techniques. Machine learning methods, with their diverse and well-designed neural network structures, along with automatic differentiation techniques, have demonstrated the ability to efficiently approximate complex functions within a vast functional space. Given this capability, it raises the question of whether we can leverage these emerging techniques to address the challenging problem of finding effective coarse-graining strategies or optimal decomposition methods for complex systems using time series data?
The answer is yes. We will introduce two latest works on causal emergence identification from data based on neural networks. Interestingly, these works indicate that causal emergence identification problem can also provide us with insight into connections with another related field, causal representation (machine) learning although causal emergence identification and causal representation learning are totally different research areas for distinct problems from appearance, they share commonalities. Furthermore, we also review the works on causal model abstractions and reinforcement learning based on world models as extensions because similar machine learning architecture are shared by these machine learning frameworks.
4.1 Causal Emergence Identification
In the previous section we have introduced several works on quantifying emergence by causality and other information theoretic conceptions. Among which, all works try to propose quantitative measures, conceptual frameworks, and numeric examples based on Markovian dynamics for causal emergence. However, in practical applications, in order for the theoretical framework to be implemented, we need to automatically identify causal emergence from real data, especially time series data of dynamic systems, and give explanations for the results.
4.1.1 Rosas’ Identification Method
The first method for causal emergence identification is given by Rosas’ paper in causal emergence Rosas et al. (2020). Actually, the three indicators, the Equations 38, 39 and 40 as mentioned in section 3.8.4. The identification criteria have been exemplified in three case studies, leading to the following conclusions: particle collisions emerge as a distinctive feature within Conway’s Game of Life, flock dynamics emerge as a characteristic feature in simulated bird behavior, and the representation of motor behavior in the cortex emerges from neural activity.
Although these three indicators avoid the problem of redundant information calculation, they are all sufficient conditions, that is, indicators greater than 0 can indicate emergence. The construction of this indicator repeatedly cuts out redundant information, which makes it almost impossible to identify the emergence of this indicator in systems with large redundant information or systems with many variables, and this is the case for most of the systems we face. Another defect of this method is that the coarse-grained variable $V$ must be given in prior, and different $V$ affects the results in a large degree. Therefore, the automatic coarse-graining strategy design according to data is still remain unsolved. New methods are required.
Therefore, the current theoretical frameworks of causal emergence do not give a good workable identification algorithm. Although existing works have tried to propose some workable methods based on static network structures and approximation for information decomposition, a general method for general Markovian dynamic systems is still lacking. One of the important reasons is to identify causal emergence within data needs to search for all the functions of coarse-graining or decomposing subsets for a complex system. However, the conventional numeric methods can not afford the computational cost for such search in a huge functional space. New methods are required.
@@ 4.1.2 Neural Information Squeezer (NIS)
Here we introduce one of the latest works on this direction which is just published in this special issue. In this paper, the authors formulate the problem of causal emergence identification on continuous Markov dynamics. Suppose the time series data $\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{T}$ are generated by a stochastic dynamical system continuously, which can be described by a differential equation,
@@
Where $\xi$ is a Gaussian noise. The time series $\{\boldsymbol{x}_{t}\}$ are the observations of the micro-states of the system, and the problem that we confront is identifying if the causal emergence occurs for the original system. And we also want to know the appropriate coarse-graining strategies, the macro dynamics which correspond to the coarse-graining method, and the dimension of the phase space on macro-level if the causal emergence occurs.
To solve the problem, the authors propose a mathematical framework by converting the original issue into an optimization problem for functions. That is, we can find the appropriate coarse-graining strategy $\phi_{q}$, which is defined on $\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$, where, $q<p$ is the dimension of the macro-states, and the macro-level Markov dynamics $f$, which is defined on $\mathcal{R}^{q}\rightarrow\mathcal{R}^{q}$ with the random noise $\zeta$, by maximizing $\mathcal{J}$, the dimension averaged EI of the macro-dynamics (see Equation 24), that is,
@@
However, this problem has a trivial solution, e.g. $\phi$ is a constant map for all micro-states, and the macro-dynamic is an identity. Surprisingly, this solution has a large EI because the identity map has the largest determinism and lowest degeneracy within all $q$ dimensional functions. This is trivial because too much information has been abandoned by the coarse-graining strategy such that the macro-dynamic is useless.
To conquer the problem, the authors add a reasonable constraint to the original optimization framework, which is called effectiveness requirement. A coarse-graining strategy and the corresponding macro-dynamic are effective, which means there exists another decoarse-graining function $\phi^{\dagger}$ such that the functions $\phi$, $f$, and $\phi^{\dagger}$ can predict the micro-state instead of the macro-state in the next time step according to the state in the previous step, the constraint can be written with,
@@
for all time step $t$, where, $\epsilon$ is a given constant. In this way, an effective coarse-graining strategy and the macro-dynamic can not only maximize the EI but also reproduce the original micro-dynamic as much as it can.
This constraint is an important complimentary for the framework of EI maximization because the decoarse-graining function $\phi^{{\dagger}}$ can map macro-states back to micro-states which can solve the problem of ambiguity of macro-states. This constraint is also a guarantee of commutativity. To see this, we can apply $(\phi^{{\dagger}})^{-1}\approx\phi$ on both sides of Equation 43 to obtain $||f(\phi(x_{t}))-\phi(x_{t+1})||<\phi(\epsilon)$.
However, the problem is still hard to solve. The paper Zhang and Liu (2022) proposes a two-stage method, which can minimize the prediction error under a given dimension $q$ of macro-states in the first step, and maximize EI for different $q$ in the second step. The first step can be solved by training a neural network easily, and the second step converts the complex functional optimization problem into a line search in one dimensional space.
Figure 14: The architecture and workflow of the Neural Information Squeezer (NIS). NIS consists of three main components: an encoder $\phi$, a dynamics learner $f$, and a decoder $\phi^{{\dagger}}$. The encoder $\phi$ comprises an invertible neural network $\psi$ and a projection operator $\chi$, which simply discards fixed dimensions output by $\phi$. On the other hand, the decoder $\phi^{{\dagger}}$ incorporates the same invertible neural network as the encoder but implements its inverse $\psi^{-1}$. In cases where the input is incomplete, it can be complemented using Gaussian random noise.
To realize the abstract mathematical framework, the authors propose an encoder-dynamics learner-decoder framework of neural network, named neural information squeezer (NIS), which describes the basic working principle of the framework vividly. To reduce the complexity of the encoder-decoder which corresponds to $\phi$ and $\phi^{\dagger}$, respectively, the authors use an invertible neural network to implement the encoder, and invert the input-output workflow of the same neural network to realize the computation of the decoder. Concretely, the encoding process is decomposed into two step computation, information conversion with the invertible neural network $\psi:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ on $p$ dimensional input space, and information discarding with the simple projection operation, that is:
@@
where $\chi_{q}$ is the projection operator to project a $p>q$ dimensional vector into a $q$ dimensional vector with the dimensions from $1$ to $q$ left. The decoding process also contains two stages, random sampling a Gaussian noise to complete the details of macro states, and information conversion but in an inverted way, that is:
@@
where, $\zeta\sim\mathcal{N}(0,I_{p-q})$ is a $p-q$ dimensional standard Gaussian random vector, and $\bigoplus$ is the operation of vector concatenation.

### 25 Feb 2024 v3

@@ 3.6.4 Reconciling Emergences
Building upon the concept of synergistic information in the PID framework, Rosas introduced a quantitative definition of causal emergence using the $\phi$ID framework to tackle the challenge of identifying an appropriate coarse-graining strategy. The definition includes two aspects: firstly, determining whether the system has the capacity to generate causal emergence, and secondly, assessing the occurrence of causal emergence given a specific macroscopic feature.
Regarding a system’s capacity to exhibit causal emergence, this definition establishes a connection between causal emergence and synergistic relationships among variables across different time points. Consequently, a system denoted as $X_{t}$ is said to possess the capacity for causally emergent features if and only if:
@@
In this context, causal emergence is understood as the synergistic effect between variables at preceding and subsequent moments within a Markovian dynamics system.
Then, Rosas further divided causal emergence into two parts in the $\phi$ID framework, downward causality and causal decoupling, based on the distinct characteristics of the information atoms. Among the sixteen $\phi$ID atoms obtained by decomposing the mutual information $I(X_{t};X_{t+1})$ using $\phi$ID, there are four information atoms corresponding to the synergistic effect, which is regarded as the composition of causal emergence. These atoms are denoted as $I^{\{12\}\rightarrow\alpha}_{\partial}(X_{t},X_{t+1})$, $\alpha\in\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$.
Downward causality is denoted as $\mathcal{D}(X_{t},X_{t+1})$, encapsulating the information atoms that exclusively manifest a synergistic effect in forward PID, as shown in Figure 1213:
$\mathcal{D}(X_{t},X_{t+1}):=\sum_{\begin{subarray}{c}\alpha\in A/\{12\}\end{subarray}}{I^{\{12\}\rightarrow\alpha}_{\partial}}(X_{t},X_{t+1})$ (35)
While causal decoupling is denoted as $\mathcal{G}(X_{t},X_{t+1})$, pinpointing the specific information atom where both forward and backward PID exhibit synergy:
@@
Since $\phi$ID is a non-overlapping decomposition of all information, this classification takes into account all cases of causal emergence, that is,
@@
Figure 12: Causal decoupling ($\mathcal{G}$) and downward causation ($\mathcal{D}$) are represented by a double-redundancy lattice. Distinct colors of dots are employed to distinguish the source elements of $X_{t}$, (colors are corresponding to Figure 11).
In addition to that, Rosas also provided an approach to quantify the causal emergence of a specific macro-variable, i.e., a coarse-graining strategy. If a system has the capacity to generate causal emergence, there can be some macroscopic features that exhibit causal emergence. A feature variable $V$ is said to be supervenient on the underlying system if it does not provide any predictive power for future states at times $t+1$ when the complete state $X$ of the system at time $t$ is known with perfect precision. This is equivalent to $V_{t}$ being statistically independent of $X_{t+1}$ given $X_{t}$. Then, for a system described by $X_{t}$, a supervenient feature $V_{t}$ is said to exhibit causal causation if:
@@
For this definition, the system’s capacity of causal emergence is required, where $Syn(X_{t};X_{t+1})$ > 0, since $Un(V_{t};X_{t+1}|X_{t})\leq Syn(X_{t};X_{t+1})$ holds for any supervenient feature $V_{t}$. Corresponding to the classification of the system’s capacity, the downward causation (as indicated by the red dotted line in Figure 13) of a feature variable $V$ exists when
@@
and causal decoupling (as indicated by the blue dotted line in Figure 13) exists when
@@
which also depends on the capacity of the system. Furthermore, $V_{t}$ is said to have pure causal decoupling if $Un(V_{t};X_{t+1}^{\alpha}|X_{t})=0$ and $Un(V_{t};X_{t+1}|X_{t})>0$. If all the emergent features exhibit pure causal decoupling, the system is said to be perfectly decoupled.
@@
Although a rigorous quantitative definition for causal emergence was proposed, the mathematical formulations used in $\phi$ID can be complex and computationally demanding, making it difficult to apply the method to real-world systems. Also, In addition, the inconsistency in PID calculation causes the definition of causal emergence to depend on a specific PID calculation. To address these problems, Rosas relaxed the calculation of causal emergence and established a identification criteria based one the sufficient conditions of causal decoupling and downward causation.
Specifically, to avoid exploring the specific quantitative methods of synergistic and redundant information, this criteria repeatedly subtracts redundant information, making the result a sufficient condition for causal emergence, which lose some of its generality but improves the reliability. The three indicators to be used are:
@@
which measures the mutual information between macro-variables across two time steps minus the mutual information between each micro-state and macro-state.
@@
which is the maximum of the difference between mutual information of $V_{t}$ and $X^{j}_{t+1}$ and the summation of the mutual information of $X_{t}^{i}$ and $X_{t+1}^{j}$.
@@
which is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$. For the above indicators, $V$ is a predefined macro-variable, but the specific method for finding such a macro-variable is not discussed in Rosas et al. (2020).
@@
The specific usage of the indicators are as follows: Firstly, $\Psi_{t,t+1}(V)>0$ is a sufficient condition for the causal emergence of $V_{t}$. Secondly, $\Delta_{t,t+1}(V)>0$ is a sufficient condition for $V_{t}$ to show downward causation. Thirdly, $\Psi_{t,t+1}(V)>0$ and $\Gamma_{t,t+1}(V)=0$ together constitute a sufficient condition for causal decoupling.
@@
In summary, Rosas proposed an approach to quantitatively characterize and classified causal emergence based on $\phi$ID by establishing the relationship between causal emergence and the synergistic effect of variables at different time points and further categorized causal emergence. The definition not only provides an objective assessment of a system’s capacity for causal emergence but also enables the measurement the causal emergence associated with a specific macro-feature. His significant contributions include bridging the gap between the study of causal emergence and quantitative empirical research, classifying the different types of causal emergence, and complementing philosophical discussions on the topic.
The PID and $\phi$ID frameworks hold the potential for explaining data in various applications. For instance, Luppi et al. recently employed the $\phi$ID method to analyze brain BOLD signals Luppi et al. (2022), aiming to identify the synergistic core of the brain. Their findings revealed that synergistic information facilitates the integration of different brain regions, whereas redundancy contributes to robustness. This study offers valuable insights into understanding the underlying mechanisms of the brain.
The advancement of partial information decomposition techniques allows for further analysis of the mutual information between variables, enabling a deeper understanding of system properties from multiple perspectives. In a study by Varley et al. Varley and Hoel (2022), the authors applied partial information decomposition to decompose the mutual information of a system. They calculated an indicator of synergy bias to assess how synergistic information is distributed across different levels of the system using the method proposed by Williams and Beer Williams and Beer (2010). A higher synergy bias indicated a greater amount of partial information involved in synergistic relationships. Subsequently, the authors observed that in certain systems exhibiting causal emergence, when the system is simplified or reduced, the synergy bias increases. This suggests that as we coarse-grain the system, partial information undergoes a transformation from redundancy to synergy. The overall conclusion drawn was that emergence can be understood as a form of information conversion.
3.6.5 Causal Emergence Identification from Data
In the previous sections, we introduced several works on quantifying emergence through causality and other information-theoretic concepts. All of these works tried to propose quantitative measures, conceptual frameworks, and numeric examples based on Markovian dynamics for causal emergence. However, in practical applications, for a theoretical framework to be implemented, we need to automatically identify causal emergence from real data, especially time-series data of dynamic systems, and provide explanations for the results.
The first method for causal emergence identification was introduced in Rosas’s paper on causal emergence Rosas et al. (2020) and involves the three indicators defined by Equations (41)–(43), as mentioned in Section 3.6.4. The identification criteria were exemplified in three case studies, leading to the following conclusions: particle collisions emerge as a distinctive feature within Conway’s Game of Life, flock dynamics emerge as a characteristic feature in simulated bird behavior, and the representation of motor behavior in the cortex emerges from neural activity.
Although these three indicators avoid the problem of redundant information calculation, it is important to note that they serve as sufficient conditions, rather than as definitive proof of emergence. In other words, indicators greater than 0 can suggest the presence of emergence, but indicators less than 0 do not necessarily imply the absence of emergence. The construction of this indicator faces challenges in identifying emergence in systems with significant redundant information or a large number of variables, which is often the case in many real-world systems. Additionally, a limitation of this method is the requirement of a predetermined coarse-grained variable $V$, and different choices of this variable can significantly impact the results. As a result, the development of an automatic coarse-graining strategy based on data remains an unresolved issue.
As a result, the current theoretical frameworks for causal emergence lack a practical and effective identification algorithm. Although previous studies have proposed methods based on static network structures and approximations for information decomposition, there is still a need for a comprehensive approach that can be applied to general Markovian dynamic systems. One of the main challenges is the necessity to search for all possible functions of coarse-graining or decomposing subsets within the data to identify causal emergence in complex systems. However, conventional numerical methods are unable to handle the computational costs associated with such an extensive search in a vast functional space. Therefore, the development of new methods is imperative to address these issues.
In Section 4, we explore the application of machine learning techniques to address the challenge of identifying causal emergence within time-series data.
@@ 3.6.6 Comparison with Hoel’s Framework
When comparing Hoel’s framework to Rosas’s quantification framework for causal emergence, several clear advantages can be observed in the latter. Firstly, Rosas’s theory does not require a predetermined coarse-graining method, making it more mathematically rigorous and formal. Secondly, it offers a detailed decomposition of causal emergence, specifically downward causation and causal decoupling. Lastly, it effectively avoids cases of fake causal emergence, where the macro-variable depends solely on unique or redundant information from the micro-variables.
@@
However, there are also some disadvantages to consider. Firstly, in order to obtain the full information lattice, a systematic iteration of all variable compositions is required. Additionally, despite the use of Formula (38), it is necessary to define a macro-variable. Unfortunately, the authors do not provide any method to identify such a variable. Secondly, all the mutual information and its decomposition are based on correlations rather than causality. It is crucial to discuss how to incorporate causal elements, such as intervention and counterfactuals, into the framework.
@@
Finally, the issue of identifying whether causal emergence occurs in a system based on the given time-series data of its behaviors has not been addressed in the preceding discussion. To address this problem, the application of emerging technologies in machine learning and artificial intelligence is required. These technologies can provide valuable tools and techniques for detecting and analyzing causal emergence.
4 Causal Emergence and Machine Learning
@@
Recently, newly emerging machine learning technologies have made significant breakthroughs in addressing a range of important and challenging problems Silver et al. (2016); Dabney et al. (2020); Senior et al. (2020); LeCun et al. (2015); Wang et al. (2018); Mikolov et al. (2013); Devlin et al. (2018); Vaswani et al. (2017); Mnih et al. (2015). Examples include defeating human champions in complex games like Go Silver et al. (2016); Dabney et al. (2020), predicting the intricate structures of protein folding Senior et al. (2020), and generating human language using large language models Fan et al. (2023). These achievements have been made possible through the application of machine learning methods, which leverage well-designed neural network architectures and automatic differentiation techniques.
However, there is a crucial limitation of conventional machine learning: it can only capture the information or associations within the data, without being able to uncover the underlying causal relationships. According to the theory of causal hierarchy by Judea Pearl Pearl et al. (2000); Pearl (2009); Cui and Athey (2022), causality is very different from association because the former always represents a more stable relationship that is invariant in different environments, whereas the latter may be more dependent on contexts and reflects the limitation of the data. Therefore, it is necessary to develop new machine learning frameworks to incorporate the consideration of causalitySchölkopf et al. (2021); Kaddour et al. (2022); Cui and Athey (2022). Numerous studies have provided evidence to show that incorporating aspects of causality can lead to improved performance in machine learning tasks, including out-of-distribution generalization, adaptation to diverse environments, and handling interventions Peters et al. (2016); Cui and Athey (2022); Kuang et al. (2018).
Machine learning and causal inference are also connected to causal emergence in two aspects, which we call “causal emergence with machine learning” and “causal emergence for machine learning”. On one hand, machine learning can be employed to address the challenge of identifying causal emergence from data, referred to as “causal emergence with machine learning”. On the other hand, causal emergence theory and effective information (EI) measures have potential applications in the field of machine learning, known as “causal emergence for machine learning”. However, the extension depends on a deeper understanding of how causal emergence theory is related to causality.
4.1 Causal Emergence with Machine Learning
In Section 3.6.5, we discussed the challenge of identifying causal emergence. Although Rosas et al. presented a method to address this issue by calculating an approximate measure of causal emergence based on their theoretical framework (Equation (41)), the method requires the availability of a macro-variable, denoted as $V_{t}$, which needs to be defined in advance. However, defining this variable in a general sense when confronted with data remains challenging. Hence, the question arises of whether we can leverage machine learning to automatically learn the macro-variable. In this section, we explore two recent approaches that utilize machine learning and neural network techniques to address the problem of causal emergence identification from the perspective of Heol et al.’s theoretical framework for EI maximization. These methods aim to learn the coarse-graining strategy, macro-variable, and macro-dynamics directly from the data, eliminating the need for explicit definitions. By leveraging the power of machine learning, we can potentially overcome the limitations of manual variable definition and enable a more automated and data-driven approach to identifying causal emergence.
@@ 4.1.1 Neural Information Squeezer (NIS)
The first work we introduce has just been published in this special issue. In this paper, the authors formulated the problem of causal emergence identification in continuous Markov dynamics. Suppose the time-series data $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$ are generated by a stochastic dynamical system continuously, which can be described by a differential equation,
@@
where $\xi$ is a Gaussian noise. The time series $\{\boldsymbol{x}_{t}\}$ represents the observations of the micro-states of the system. The problem we confront is identifying whether causal emergence occurs for the original system. We also want to know the appropriate coarse-graining strategies, the macro-dynamics corresponding to the chosen coarse-graining method, and the dimension of the phase space at the macro-level when the causal emergence occurs.
To address this problem, the authors proposed a mathematical framework by converting the original issue into an optimization problem for functions. Specifically, the goal is to find the appropriate coarse-graining strategy $\phi_{q}$, defined on $\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$, where $q<p$ is the dimension of the macro-states. Additionally, the aim is to determine the macro-level Markov dynamics $f$, defined on $\mathcal{R}^{q}\rightarrow\mathcal{R}^{q}$ with the random noise $\zeta$, by maximizing $\mathcal{J}$—the dimension-averaged EI of the macro-dynamics (see Equation (27)). This is expressed as,
@@
However, this problem has a trivial solution, e.g., $\phi$ is a constant map for all micro-states, and the macro-dynamic is an identity. Surprisingly, this solution has a large EI because the identity map has the largest determinism and lowest degeneracy among all $q$-dimensional functions. This is trivial because too much information has been abandoned by the coarse-graining strategy such that the macro-dynamic is useless.
To address the problem, the authors added a reasonable constraint to the original optimization framework, called the effectiveness requirement. A coarse-graining strategy and its corresponding macro-dynamics are considered effective if there exists another decoarse-graining function $\phi^{\dagger}$ such that the functions $\phi$, $f$, and $\phi^{\dagger}$ can predict the micro-state, instead of the macro-state, in the next time step based on the state in the previous step. The constraint can be written as,
@@
for all time steps $t$, where $\epsilon$ is a given constant. This way, an effective coarse-graining strategy and macro-dynamics not only maximize the EI but also reproduce the original micro-dynamics as much as possible.
This constraint is an important complement to the framework for EI maximization because the decoarse-graining function $\phi^{{\dagger}}$ can map macro-states back to micro-states, addressing the problem of the ambiguity of macro-states. This constraint also ensures commutativity. To demonstrate this, we can apply $(\phi^{{\dagger}})^{-1}\approx\phi$ to both sides of Equation (46), resulting in $||f(\phi(x_{t}))-\phi(x_{t+1})||<\phi(\epsilon)$.
However, the problem is still hard to solve. In Zhang and Liu (2022), the authors proposed a two-stage method, which can minimize the prediction error under a given dimension $q$ of macro-states in the first step and maximize EI for different values of $q$ in the second step. The first step can be solved by training a neural network, and the second step converts the complex functional optimization problem into a line search in one-dimensional space.
To realize the abstract mathematical framework, the authors proposed an encoder-dynamics learner-decoder framework for neural networks, named the Neural Information Squeezer (NIS), vividly describing the basic working principle of the framework as depicted in Figure 14.
Figure 14: The architecture and workflow of the Neural Information Squeezer (NIS). The NIS consists of three main components: an encoder $\phi$, a dynamics learner $f$, and a decoder $\phi^{{\dagger}}$. The encoder $\phi$ comprises an invertible neural network $\psi$ and a projection operator $\chi$, which simply discards fixed dimensions output by $\phi$. On the other hand, the decoder $\phi^{{\dagger}}$ incorporates the same invertible neural network as the encoder but implements its inverse $\psi^{-1}$. In cases where the input is incomplete, it can be complemented using Gaussian random noise.
In order to reduce the complexity of the encoder and decoder, corresponding to $\phi$ and $\phi^{\dagger}$, respectively, the authors used an invertible neural network to implement the encoder and inverted the input-output workflow of the same neural network to implement the decoder. Concretely, the encoding process was decomposed into two computation steps: information conversion with the invertible neural network $\psi:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ on a $p$-dimensional input space, and information discarding with a simple projection operation. This is expressed as:
@@
where $\chi_{q}$ is the projection operator that projects a $p>q$-dimensional vector into a $q$-dimensional vector, retaining dimensions from $1$ to $q$. The decoding process also consists of two stages: randomly sampling Gaussian noise to complete the details of macro-states, and information conversion but in an inverted way, that is,
@@
where $\zeta\sim\mathcal{N}(0,I_{p-q})$ is a $p-q$-dimensional standard Gaussian random vector, and $\bigoplus$ is the operation of vector concatenation.